OpenStudy (anonymous):

I've done this problem a few times and can't get it. Help? It is using quadratic form, letting U=x^2 24x^4+14x^2-3=0

OpenStudy (anonymous):

srry but she might be able 2 help @kristyn-whitaker

OpenStudy (displayerror):

Use the substitution that is provided. Your equation as given: \[24x^4 + 14x^2-3=0\] So we can replace the x terms in the first equation with U by using the substitution as follows \[U^2 = (x^2)^2 = x^4\] \[U=x^2\] The equation then becomes \[24U^2 + 14U -3 =0\] Now that is in a form that you should be able to solve by either factoring or using the quadratic equation.